# Runtime Safety through Adaptive Shielding: From Hidden Parameter Inference to Provable Guarantees

### v1

@@ 5 Experiments
We empirically evaluate our approach to assess its safety, generalization, and efficiency across diverse RL tasks  [^1] Our code is available on Code Link. We compare against established safe RL baselines and analyze three variants of our method: using only the safety-regularized objective, only the adaptive shield, and their combination. Our experiments are guided by the following research questions:
  • RQ1: How does our approach balance safety and task performance during training?
  • RQ2: How well does our approach generalize to out-of-distribution test environments?
  • RQ3: What is the runtime overhead of our approach at execution?
@@ 5.1 Experimental Setup
Environments. We conduct experiments using the Safe-Gym benchmark (Ji et al., 2023) for safe RL, with two robot types: Point and Car. Each robot performs four tasks: (1) Goal: navigate to a target while avoiding obstacles; (2) Button: activate a button while avoiding hazards; (3) Push: push an object to a goal under contact constraints; (4) Circle: follow a circular path while staying within safe boundaries. Robot-task combinations are denoted as robot-task (e.g., Point-Goal, Car-Circle). Each task includes a safety constraint (e.g., obstacle avoidance or region adherence). Episode-level randomness is introduced by sampling gravity ($\alpha_{g}$) and hidden dynamics parameters ($\alpha_{d}$, $\alpha_{\rho}$).
Baselines. We compare our approach to four established safe RL algorithms:
  • TRPO-Lag: Trust Region Policy Optimization with second-order reward updates and first-order constraint updates (Schulman et al., 2015).
@@
  • PPO-Lag: Proximal Policy Optimization with Lagrangian updates for both reward and constraint (Schulman et al., 2017).
@@
  • CPO: Constrained Policy Optimization with joint second-order updates to enforce linearized cost constraints (Achiam et al., 2017).
@@
  • USL: Unrolling Safety Layer, which re-weights the policy loss for safety and projects unsafe actions into a feasible set at execution (Zhang et al., 2023).
Baselines are given access to hidden parameters to support dynamics adaptation, as they do not perform inference. In contrast, our approach uses $\hat{f}_{\text{FE}}$ to infer hidden parameters online and is evaluated without this privileged information, demonstrating robustness under limited parameter awareness.
@@
Hyperparameters. All methods use the default hyperparameters provided by their respective implementations: Omni-Safe (Ji et al., 2024) for TRPO-Lag, PPO-Lag, and CPO, and Safe-RL-Kit (Zhang et al., 2023) for USL. When evaluating our approach on top of each base algorithm (e.g., TRPO or PPO), we adopt the same hyperparameters as the corresponding baseline to ensure a fair comparison. Each method is trained for 2 million environment steps using 3 random seeds. Each trained policy is evaluated over 100 episodes at test time.
@@
We set the pre-safety distance to 0.275 and the ACP failure probability to 2%. The function encoder $\hat{f}_{\text{FE}}$ is pre-trained on 1000 episodes (1000 steps each) and remains fixed during policy training, introducing realistic prediction error that is managed by ACP. All agents are trained under a strict safety constraint, with a cost limit of zero.
@@
For in-distribution training, environment parameters ($\alpha_{g}$, $\alpha_{d}$, $\alpha_{\rho}$) are sampled uniformly from $[0.25,1.75]$. For out-of-distribution evaluation, parameters are sampled from $[0.1,0.25]\cup[1.75,2.5]$, and the number of obstacles is increased to stress generalization.
@@
Metrics. We evaluate each method using per-episode averages for the following metrics, each capturing a different aspect of performance: (1) Return, measuring task performance as the cumulative reward per episode; (2) Cost Rate, reflecting safety by measuring the frequency of constraint violations per timestep; (3) Runtime, quantifying execution-time efficiency as the wall-clock time per episode; and (4) Shielding Rate, indicating how often the shield is triggered during an episode (i.e., when the pre-safety check fails).
@@ 5.2 Results Analysis
Figure 1: Training curves of episodic return and cost rate for the Car robot across four tasks. Solid lines show means over three seeds; shaded regions indicate the 95% range as $\pm$2 standard deviations.
RQ1: Trade-offs Between Safety and Return. Figure 1 shows training curves of episodic return and cost rate for the Car robot across four tasks. Baseline methods exhibit a range of trade-offs. TRPO-Lag and PPO-Lag tend to achieve high returns but incur higher cost rates. CPO enforces strict safety through a zero-violation constraint, often at the expense of reward learning, resulting in ineffective policies on several tasks. USL, which relies on cost-Q-value estimation, performs well on the Circle task but struggles on Goal, Button, and Push—likely due to sensitivity to environment stochasticity, such as randomly reset obstacle positions, which disrupt cost estimation. These results highlight the challenge of achieving both safety and return without mechanisms for adapting to latent dynamics and managing uncertainty.
In contrast, our methods (using TRPO as the base RL algorithm) consistently achieve lower cost rates while maintaining competitive returns, demonstrating their ability to balance safety and task performance during training. Variants using only the safety-regularized objective (SRO) or only the adaptive shield also reduce cost violations compared to baselines, but are less effective than the combined method. We observe similar results for the Point robot and when using PPO as the base RL algorithm for our methods (see results in Appendix C and D).
Takeaway: our combined method (SRO + shield) achieves the best safety outcomes during training, with only a modest impact on task return.
Figure 2: Trade-off between average episodic return and cost rate in out-of-distribution domains.
@@
Our full method (SRO + Shield) consistently appears near or within the desirable yellow region (high return and low cost rate) across all tasks, indicating strong generalization to previously unseen dynamics. The SRO-only and Shield-only variants also perform well but tend to deviate more from the Pareto front, especially in complex environments. This trend highlights the complementary effect of combining proactive (SRO) and reactive (Shield) safety mechanisms.
Among the baselines, CPO maintains low cost rates but sacrifices return, often placing it outside the desirable region. PPO-Lag and TRPO-Lag yield high returns but suffer from high violation rates. USL exhibits inconsistent behavior, particularly on the Point robot, suggesting poor generalization under dynamic, partially-observed conditions.
Takeaway: our approach generalizes effectively to out-of-distribution settings, consistently achieving a favorable balance between return and safety across robot types and task variations.
RQ3: Execution-Time Efficiency. Table 1 compares the average runtime per episode between the baseline method (TRPO-Lag) and our full approach (SRO + Shield). Across all tasks and both robot types, our method introduces only a modest runtime overhead, demonstrating practical efficiency during execution. Additionally, the shield trigger rate remains moderate, indicating that safety shielding is invoked selectively and does not dominate execution time.
Takeaway: our approach achieves safe execution with limited computational overhead, making it suitable for real-time deployment.
@@
Car TRPO-Lag Runtime (s) 2.52$\pm$0.03 2.97$\pm$0.02 2.85$\pm$0.01 1.74$\pm$0.00
Ours Runtime (s) 2.80$\pm$0.04 3.10$\pm$0.06 3.11$\pm$0.02 1.59$\pm$0.08
Ours Shielding Rate (%) 22.15$\pm$2.88 16.68$\pm$1.80 17.05$\pm$1.52 15.82$\pm$0.29
Point TRPO-Lag Runtime (s) 2.26$\pm$0.00 2.62$\pm$0.01 2.58$\pm$0.03 1.63$\pm$0.02
Ours Runtime (s) 2.57$\pm$0.15 2.70$\pm$0.00 2.69$\pm$0.06 1.71$\pm$0.07
Ours Shielding Rate (%) 22.55$\pm$5.23 14.09$\pm$0.50 13.66$\pm$3.00 5.93$\pm$5.23
@@
Ablation Studies. We conduct additional ablation studies, presented in the appendix. These include an analysis of key hyperparameters, such as sampling size and safety bonus (Appendix E), and an evaluation of the impact of the function encoder’s representation on safety and performance (Appendix F). Our ablations show that performance remains stable across different sampling sizes and safety bonus values. Additionally, our function encoder-based inference effectively adapts to varying dynamics $T$, achieving performance comparable to that of oracle representations.
@@ 6 Conclusion
We presented a novel approach for safe and generalizable reinforcement learning in settings with dynamically varying hidden parameters. Our approach comprises three key components: (1) function encoder-based inference of hidden dynamics, (2) a safety-regularized objective that promotes low-violation behavior during training, and (3) an adaptive runtime shield that uses conformal prediction to filter unsafe actions based on uncertainty at execution time. Experimental results demonstrate that our approach consistently outperforms strong baselines in reducing safety violations while maintaining competitive task performance, and generalizes effectively across diverse tasks and out-of-distribution environments.
@@
Appendix A Broader Impacts
This paper discusses potential positive societal impacts of our safe RL framework for robotics. We do not identify specific negative societal impacts from our research, as it focuses on improving safety and robustness in controlled settings. Our work enhances safety in autonomous systems, which could benefit applications like autonomous driving and robotic manipulation by reducing accidents.
Appendix B Proofs
@@
Notations. We introduce the notation for dimensions and sets as follows. For $n,m,k_{1},k_{2}\in\mathbb{N}$, let $n$ denote the state dimension, $m$ the action dimension, $k_{1}$ the hidden parameter dimension, and $k_{2}$ the dimension of the function encoder’s learned representation. Note that the dimension $k_{2}$ of the learned representation may differ from $k_{1}$, depending on the number of chosen basis functions. The state space is $S\subseteq\mathbb{R}^{n}$, the action space is $A\subseteq\mathbb{R}^{m}$, the hidden parameter space is $\Theta\subseteq\mathbb{R}^{k_{1}}$, and the learned representation space is $\Phi\subseteq\mathbb{R}^{k_{2}}$, where $\Phi$ is the coefficient space of basis functions for the function encoder.
@@
  • $\theta\in\mathbb{R}^{k_{1}}$: hidden parameter.
  • $b\in\mathbb{R}^{k_{2}}$: learned representation to $T_{\theta}$.
  • $s^{\prime}\sim T(\cdot\mid s,a,\theta)$: Transition dynamics given parameter $\theta$. When $s,a,s^{\prime}$ are unspecified, we denote this by $T_{\theta}$.
@@
  • $\mathcal{S}_{\text{safe}}(s_{t},a_{t})$: Safe state set, defined as $\{s_{t}\in\mathcal{S}\mid C(s_{t},a_{t},s_{t+1})=0\text{ s.t }T(s_{t+1}\mid s_{t},a_{t},\theta)>0\}$ where $\theta$ is a parameter sampled per episode.
  • $Q_{R}^{\pi}(s,a,\theta)$: State-action value function for reward under policy $\pi$ defined as
    $\mathbb{E}_{\pi,T_{\theta}}\left[\sum_{t=0}^{\infty}\gamma^{t}R(s_{t},a_{t},s_{t+1})\mid s_{0}=s,a_{0}=a,\theta\right].$
@@
  • $Q^{\pi}_{C}(s,a,\theta)$: State-action value function for cost under policy $\pi$ defined as
    $\mathbb{E}_{\pi,T_{\theta}}\left[\sum_{t=0}^{\infty}\gamma^{t}C(s_{t},a_{t},s_{t+1})\mid s_{0}=s,a_{0}=a,\theta\right].$
@@
  • $Q^{\pi}_{\text{aug}}(s,a,\theta)$: Safety-regularized state-action value function defined as $Q_{R}^{\pi}(s,a,\theta)+\alpha Q_{\text{safe}}^{\pi}(s,a,\theta)$ where $\alpha$ is a positive constant and
    $Q^{\pi}_{\text{safe}}(s,a,\theta)=1-\frac{\pi(a\mid s,\theta)Q^{\pi}_{C}(s,a,\theta)}{V^{\pi}_{C}(s,\theta)+\epsilon}$
    with $\epsilon>0$ is a small constant for numerical stability. We aim to maximize this value.
  • $J_{\text{safe}}(\pi)$: Safety-regularized objective function for policy $\pi$ defined as
    $\mathbb{E}_{\theta\sim P_{\Theta},s_{0}\sim\mu_{0}(\cdot\mid\theta),a_{0}\sim\pi(\cdot\mid s_{0},b)}\left[Q^{\pi}_{\text{aug}}(s_{0},a_{0},\theta)\right].$
@@
    $\mathbb{E}_{\theta\sim P_{\Theta},s_{0}\sim\mu_{0}(\cdot\mid\theta),a_{0}\sim\pi(\cdot\mid s_{0},b)}\left[Q^{\pi}_{R}(s_{0},a_{0},\theta)\right].$
@@
    $\mathbb{E}_{\theta\sim P_{\Theta},s_{0}\sim\mu_{0}(\cdot\mid\theta),a_{0}\sim\pi(\cdot\mid s_{0},b)}\left[Q^{\pi}_{C}(s_{0},a_{0},\theta)\right].$
@@
  • $\phi^{\pi}(s,\theta)$: Average cost under policy $\pi$ defined as
    $\phi^{\pi^{*}}\left(s,\theta\right)=\lim_{H\rightarrow\infty}\frac{1}{H}\mathbb{E}_{\pi^{*},T_{\theta}}\left[\sum_{t=0}^{H-1}C(s_{t},a_{t},s_{t+1})\mid s_{0}=s,\theta\right],$
    starting from state $s$ and parameter $\theta$.
@@
Proposition 1. Let $\Pi_{\text{zero-violation}}$ be the set of zero-violation policies. Then, for any $\alpha\geq 0$, the optimal policy obtained by maximizing the safety-regularized objective function $J_{\text{safe}}(\pi)$ within $\Pi_{\text{zero-violation}}$ is equivalent to the optimal policy obtained by maximizing the standard reward objective $J_{R}(\pi)$ within the same set of policies.
@@
By definition, if $\pi\in\Pi_{\text{zero-violation}}$, then for any state $s$, parameter $\theta$, and action $a$ with $\pi(a|s,\theta)>0$, the state-action value function for the cost is zero: $Q^{\pi_{\theta}}_{C}(s,a,\theta)=0$. This implies that for such policies, the safety term $Q^{\pi}_{\text{safe}}(s,a,\theta)$ in the regularized objective is a constant value of $1$.
@@
$J_{\text{safe}}(\pi)$ $=\mathbb{E}_{\theta\sim P_{\Theta},s_{0}\sim\mu_{0}(\cdot\mid\theta),a_{0}\sim\pi(\cdot\mid s_{0},b)}\left[Q_{R}^{\pi}(s_{0},a_{0},\theta)+\alpha\cdot Q^{\pi}_{\text{safe}}(s_{0},a_{0},\theta)\right]$
$=\mathbb{E}_{\theta\sim P_{\Theta},s_{0}\sim\mu_{0}(\cdot\mid\theta),a_{0}\sim\pi(\cdot\mid s_{0},b)}\left[Q_{R}^{\pi}(s_{0},a_{0},\theta)+\alpha\cdot 1\right]$
$=J_{R}(\pi)+\alpha$
Since $\alpha$ is a constant, maximizing $J_{\text{safe}}(\pi)$ within $\Pi_{\text{zero-violation}}$ is equivalent to maximizing $J_{R}(\pi)$ within the same set. ∎
@@
We define the safe state set for a given state $s\in S$, action $a\in A$, and hidden parameter $\theta\in\Theta$ as: $\mathcal{S}_{\mathrm{safe}}(s,a)=\left\{s^{\prime}\in S\mid C(s,a,s^{\prime})=0\text{ s.t. }T\left(s^{\prime}\mid s,a,\theta\right)>0\right\}.$ Thus, $\mathcal{S}_{\text{safe}}(s,a)$ contains next states $s^{\prime}$ where the safety condition is satisfied. Throughout the theorem, we address a cost function defined by safe distance

### v2

@@ 5 Experiments
We empirically evaluate our approach to assess its safety, generalization, and efficiency across diverse RL tasks. We compare against six safe RL algorithms and analyze three variants of our method: using only safety-regularized optimization (SRO), only the adaptive shield, and their combination. Our experiments are guided by the following research questions:
  • RQ1: How does our approach balance safety and task performance during training without being informed of changing hidden parameters?
  • RQ2: How well does our approach generalize to out-of-distribution test environments by inferring varying hidden parameters online?
  • RQ3: What is the execution-time overhead introduced by the adaptive shielding mechanism?
@@ 5.1 Experimental Setup
Environments. We conduct experiments using the Safe-Gym benchmark (ji2023safety) for safe RL, with two robot types: Point and Car. Each robot performs four tasks: (1) Goal: navigate to a target while avoiding obstacles; (2) Button: activate a button while avoiding hazards; (3) Push: push an object to a goal under contact constraints; (4) Circle: follow a circular path while staying within safe boundaries. Robot-task combinations are denoted as robot-task (e.g., Point-Goal, Car-Circle). Each task includes a safety constraint (e.g., obstacle avoidance or region adherence). Episode-level randomness is introduced by sampling gravity, and four hidden dynamics parameters: damping, mass, inertia, and friction.
Baselines. We compare our approach to six safe RL algorithms:
  • Saute: A state augmentation technique with safety budgets for almost sure constraint satisfaction, applicable to a wide range of RL algorithms such as PPO or RCPO (li2022saute).
@@
  • PPO-Lag: Proximal Policy Optimization with Lagrangian updates for both reward and constraint (schulman2017ppo; ray2019benchmarking).
  • RCPO: Reward Constrained Policy Optimization, which uses policy gradients to optimize a reward function penalized by safety violations (Chen2018).
@@
  • CPO: Constrained Policy Optimization with joint second-order updates to enforce linearized cost constraints (achiam2017constrained).
  • CUP: Constrained Update Projection, a policy optimization method that projects updates to satisfy safety constraints with theoretical guarantees (yang2022cup).
@@
  • USL: Unrolling Safety Layer, which re-weights the policy loss for safety and projects unsafe actions into a feasible set at execution (zhang2023evaluating).
We compare against oracle baselines that receive the ground-truth hidden parameters $\phi$ as input. This represents a theoretical upper bound for any history-based inference method (e.g., Transformer), which can at best recover $\phi$. In contrast, our approach uses $\hat{f}_{\text{FE}}$ to infer hidden parameters online and is evaluated without this oracle information, demonstrating robustness under limited parameter awareness. We use RCPO without access to the hidden parameters $\phi$ as the base RL algorithm. On top of this, we evaluate RCPO combined with SRO, RCPO combined with Shield, and RCPO combined with both SRO and Shield. For brevity, we refer to these as SRO, Shield, and SRO + Shield, respectively. We also provide results using PPO-Lag combined with our methods in Appendix F.
@@
Hyperparameters. All methods use the default hyperparameters provided by their respective implementations: Omni-Safe (omnisafe) for Saute, PPO-Lag, RCPO, CPO, and CUP, and Safe-RL-Kit (zhang2023evaluating) for USL. When evaluating our approach on top of each base algorithm (e.g., RCPO), we adopt the same hyperparameters as the corresponding baseline to ensure a fair comparison. Each method is trained for 2 million environment steps using 3 random seeds. Each trained policy is evaluated over 100 episodes at test time.
@@
We set the pre-safety distance to 0.275 and the ACP failure probability to 2%. The function encoder $\hat{f}_{\text{FE}}$ is pre-trained on 1000 episodes (1000 steps each) collected by a trained PPO policy. The $\hat{f}_{\text{FE}}$ remains fixed during policy training, introducing realistic prediction error that is managed by ACP. All agents are trained under a strict safety constraint, with a cost limit of zero.
@@
For training, environment parameters $\phi$ (gravity, damping, mass, inertia, friction) are sampled uniformly from the interval $[0.3,1.7]$. For out-of-distribution evaluation, the parameters are sampled from the interval $0.15,0.3]\cup[1.7,2.5]$, and the number of obstacles is increased to stress generalization.
@@
Metrics. We evaluate each method using per-episode averages for the following metrics, each capturing a different aspect of performance: (1) Return, measuring task performance as the cumulative reward per episode; (2) Cost Rate, reflecting safety by measuring the frequency of constraint violations per timestep. (3) Runtime, quantifying execution-time efficiency as the wall-clock time per episode; and (4) Shielding Rate, indicating how often the shield is triggered during an episode (i.e., when the pre-safety check fails).
@@ 5.2 Results Analysis
Figure 1: Training Dynamics. Results display the mean reward and cost rate (%) over the last 20 epochs across seeds. The top-left position is desirable, indicating higher returns with lower cost rates. Solid points represent mean return and cost rate, while transparent points depict individual seed results.
RQ1: Trade-offs Between Safety and Return. Figure 1 shows the episodic return and cost rate across four tasks during training. Baseline methods exhibit a range of trade-offs. PPO-Lag tends to achieve high returns but incur higher cost rates. CPO, PPO-Saute, and RCPO-Saute enforce strict safety via a zero-violation constraint, often sacrificing reward learning, which leads to suboptimal policies in multiple tasks. USL, dependent on cost-Q-value estimation, underperforms across all tasks, particularly in environments with randomized obstacle configurations. CUP preserves task-solving performance but frequently violates strict cost limits. RCPO maintains a reasonable balance between safety and return but struggles to meet lower cost thresholds. These results highlight the challenge of balancing safety and returns, even when hidden parameters are provided as inputs, in the presence of varying hidden parameters. To assess the impact of providing fixed parameters, we compare performance with and without fixed parameters. See Appendix K for details.
In contrast, our methods (using RCPO as the base RL algorithm) consistently achieve lower cost rates while maintaining competitive returns, demonstrating their ability to balance safety and task performance during training. Variants using only SRO or only adaptive shield also reduce cost violations compared to baselines, but are less effective than the combined method. We observe similar results when using PPO-Lag as the base RL algorithm for our methods (see Appendix F).
Takeaway: The combined method (SRO + Shield) achieves a favorable reward-cost trade-off during training and maintains stable performance under unseen variations in environmental parameters.
Figure 2: OOD Evaluation. Trade-off between average episodic return and cost rate in out-of-distribution domains. The top-left position is desirable, indicating higher returns with lower cost rates. Solid points represent mean return and cost rate, while transparent points depict individual seed results.
@@
Our combined method (SRO + Shield) consistently appears near the desirable position (high return and low cost rate) across all tasks, indicating robust adaptation to previously unseen dynamics. SRO-only and Shield-only variants also perform well but tend to deviate more from the desirable position. Shield typically remains stable because its logic, which filters unsafe actions based on predicted states and safety measures like proximity to obstacles, holds regardless of OOD conditions. This trend highlights the complementary effect of combining proactive (SRO) and reactive (Shield) safety mechanisms.
Among the baselines, CPO and Saute maintain low cost rates but sacrifice return, often positioning them outside the desirable region. The remaining algorithms exhibit high variance and sensitivity to dynamics shifts, failing to maintain a consistent safety-performance trade-off across environments.
Takeaway: The combined method effectively bridges the gap between safety and adaptation, maintaining a favorable return-safety trade-off under out-of-distribution dynamics where non-adaptive baselines fail.
RQ3: Execution-Time Efficiency. Table 1 compares the average runtime per episode between the baseline method (RCPO) and our combined approach (SRO + Shield). Across all tasks and both robot types, the proposed method introduces a modest execution-time overhead relative to the base policy. Additionally, the shield trigger rate remains moderate, indicating that safety shielding is invoked selectively and does not dominate execution time.
Takeaway: The combined method incurs limited additional computational cost over the base policy while providing consistent runtime safety filtering.
@@
[…]
@@
Ablation Studies. We provide additional analyses in the Appendix to assess the robustness of our proposed shielding framework. Key findings include: (1) Compatibility to other Algorithms: We demonstrate that the proposed runtime shielding mechanism improves safety across different base RL algorithms (Appendix F). This demonstrate that our adaptive shielding does not depend on a specific base algorithm. (2) Dynamics Representation: We show that Function Encoder shows more stable dynamics representations than a Transformer-based alternative, and achieves next-state prediction accuracy comparable to oracle-informed predictors under distribution shift (Appendices G and H). (3) Transfer to Higher-Dimensional Control: We extend evaluation to HalfCheetah-Velocity task with multiple hidden parameters, providing evidence that the proposed approach performs well in more complex task environments (Appendix I). Additional sensitivity analyses on safety-bonus weight and action-sampling size are shown in Appendix E. We derive how SRO in theory connects to TRPO or CPO algorithms in Appendix B. We provide supporting evidence for our SRO design choice in Appendix J. We provide the relationship between our proposed adaptive shield and classical control theory in Appendix D.
@@ 6 Conclusion
We present an adaptive shielding framework for safe and generalizable reinforcement learning that addresses the challenge of maintaining safety while adapting to varying hidden dynamics. By using function encoders, the shield performs inference of hidden dynamics. During training, safety-regularized optimization promotes low-violation behavior in the policy. At execution time, the adaptive shield uses conformal prediction to filter unsafe actions based on uncertainty. Across out-of-distribution Safe-Gym benchmarks, we empirically demonstrate consistent reductions in safety violations relative to baselines while preserving competitive task performance and incurring only modest computational overhead. Our theoretical analysis characterizes how prediction errors in the learned dynamics model connect with bounds on the average cost rate of the shielded policy.
@@
Reproducibility Statement
To ensure the reproducibility of our results, we provided the complete source code including a Bash script to execute the entire training and evaluation pipeline. This streamlined approach facilitates use by the research community and provide a reliable baseline for Safe and generalizable RL.
Broader Impacts
This work addresses how to maintain safety for RL agents under unseen distribution shifts in environment dynamics caused by hidden parameters (e.g., friction, mass, gravity). We propose an adaptive, uncertainty-aware shielding framework that infers a low-dimensional dynamics representation online and filters unsafe actions at execution time.
Benefits. Runtime shielding can reduce constraint violations without retraining, improving reliability of learning-based controllers in safety-critical domains such as robotics and autonomous systems. The use of uncertainty-aware bounds encourages conservative behavior when predictions are unreliable.
Risks. A primary risk is false confidence: probabilistic safety guarantees may be over-trusted outside our modeling assumptions or when the offline data used to train the function encoder poorly covers extreme dynamics. In such cases, safety filtering may degrade or fail.
Mitigations. We recommend treating shielding as one layer in a safety stack, and stress testing under harder distribution shifts, rare events, and adversarial parameter settings. Reporting trigger rates and failure cases can help detect operation outside the method’s reliable guarantees.
Appendix A Proofs
@@
Notations. We introduce the notation for dimensions and sets as follows. For $n,m,k_{1},k_{2}\in\mathbb{N}$, let $n$ denote the state dimension, $m$ the action dimension, $k_{1}$ the hidden parameter dimension, and $k_{2}$ the dimension of the function encoder’s learned representation. Note that the dimension $k_{2}$ of the learned representation may differ from $k_{1}$, depending on the number of chosen basis functions. The state space is $S\subseteq\mathbb{R}^{n}$, the action space is $A\subseteq\mathbb{R}^{m}$, the hidden parameter space is $\Phi\subseteq\mathbb{R}^{k_{1}}$, and the learned representation space is $\mathcal{B}\subseteq\mathbb{R}^{k_{2}}$, where $\mathcal{B}$ is the coefficient space of basis functions for the function encoder.
@@
  • $\phi\in\mathbb{R}^{k_{1}}$: hidden parameter.
  • $b_{\phi}\in\mathbb{R}^{k_{2}}$: learned representation to $T_{\phi}$.
  • $s^{\prime}\sim T(\cdot\mid s,a,\phi)$: Transition dynamics given parameter $\phi$. When $s,a,s^{\prime}$ are unspecified, we denote this by $T_{\phi}$.
@@
  • $\mathcal{S}_{\text{safe}}(s_{t},a_{t})$: Safe state set, defined as $\{s_{t}\in\mathcal{S}\mid C(s_{t},a_{t},s_{t+1})=0\text{ s.t }T(s_{t+1}\mid s_{t},a_{t},\phi)>0\}$ where $\phi$ is a parameter sampled per episode.
  • $Q_{R}^{\pi}(s,a,b_{\phi})$: State-action value function for reward under policy $\pi$ defined as
    $\mathbb{E}_{\pi,T_{\phi}}\left[\sum_{t=0}^{\infty}\gamma^{t}R(s_{t},a_{t},s_{t+1})\mid s_{0}=s,a_{0}=a,\phi\right].$
@@
  • $Q^{\pi}_{C}(s,a,b_{\phi})$: State-action value function for cost under policy $\pi$ defined as
    $\mathbb{E}_{\pi,T_{\phi}}\left[\sum_{t=0}^{\infty}\gamma^{t}C(s_{t},a_{t},s_{t+1})\mid s_{0}=s,a_{0}=a,\phi\right].$
@@
  • $Q^{\pi}_{\text{aug}}(s,a,b_{\phi})$: Safety-regularized state-action value function defined as $Q_{R}^{\pi}(s,a,b_{\phi})+\alpha Q_{\text{safe}}^{\pi}(s,a,b_{\phi})$ where $\alpha$ is a positive constant and
    $Q^{\pi}_{\text{safe}}(s,a,b_{\phi})=-\frac{\int_{B(a,\epsilon)\cap A}\pi(x\mid s,b_{\phi})Q^{\pi}_{C}(s,x,b_{\phi})dx}{V^{\pi}_{C}(s,b_{\phi})+\epsilon}$
    where $\epsilon>0$ is a small constant ensuring numerical stability, and $B(a,\epsilon)$ denotes a small ball of radius $\epsilon$ centered at $a$. We aim to maximize this value, whose maximum value is 0.
  • $J_{\text{safe}}(\pi)$: Safety-regularizer for policy $\pi$ defined as
    $\mathbb{E}_{\phi\sim P_{\phi},s_{0}\sim\mu_{0}(\cdot\mid\phi),a_{0}\sim\pi(\cdot\mid s_{0},b_{\phi})}\left[Q^{\pi}_{\text{safe}}(s_{0},a_{0},b_{\phi})\right].$
    The policy $\pi$ aims to maximize this value.
  • $J_{\text{aug}}(\pi)$: Safety-regularized objective function for policy $\pi$ defined as
    $\mathbb{E}_{\phi\sim P_{\phi},s_{0}\sim\mu_{0}(\cdot\mid\phi),a_{0}\sim\pi(\cdot\mid s_{0},b_{\phi})}\left[Q^{\pi}_{\text{aug}}(s_{0},a_{0},b_{\phi})\right].$
@@
    $\mathbb{E}_{\phi\sim P_{\phi},s_{0}\sim\mu_{0}(\cdot\mid\phi),a_{0}\sim\pi(\cdot\mid s_{0},b_{\phi})}\left[Q^{\pi}_{R}(s_{0},a_{0},b_{\phi})\right].$
@@
    $\mathbb{E}_{\phi\sim P_{\phi},s_{0}\sim\mu_{0}(\cdot\mid\phi),a_{0}\sim\pi(\cdot\mid s_{0},b_{\phi})}\left[Q^{\pi}_{C}(s_{0},a_{0},b_{\phi})\right].$
@@
  • $\xi^{\pi}(s,\phi)$: Average cost under policy $\pi$ defined as
    $\xi^{\pi^{*}}\left(s,\phi\right)=\lim_{H\rightarrow\infty}\frac{1}{H}\mathbb{E}_{\pi^{*},T_{\phi}}\left[\sum_{t=0}^{H-1}C(s_{t},a_{t},s_{t+1})\mid s,\phi\right],$
    starting from an initial state $s$ and parameter $\phi$.
@@
Proposition 1. Let $\Pi_{\text{zero-violation}}$ be the set of zero-violation policies. Then, for any $\alpha\geq 0$, the optimal policy obtained by maximizing the safety-regularized objective function $J_{\text{aug}}(\pi)$ within $\Pi_{\text{zero-violation}}$ is equivalent to the optimal policy obtained by maximizing the standard reward objective $J_{R}(\pi)$ within the same set of policies.
@@
By definition, if $\pi\in\Pi_{\text{zero-violation}}$, then for any state $s$, parameter $\phi$, and action $a$ with $\pi(a|s,b_{\phi})>0$, the state-action value function for the cost is zero: $Q^{\pi}_{C}(s,a,b_{\phi})=0$. Hence, any action sampled around $a$ by the policy $\pi$ will have $Q^{\pi}_{C}(s,a,b_{\phi})=0$ leading to
$\int_{B(a,\epsilon)\cap A}\pi(x\mid s,b_{\phi})Q^{\pi}_{C}(s,x,b_{\phi})dx=0.$
This implies that for such policies, the safety term $Q^{\pi}_{\text{safe}}(s,a,b_{\phi})$ in the regularized objective is a constant value of 0.
@@
$J_{\text{aug}}(\pi)$
$=\mathbb{E}_{\phi\sim P_{\phi},s_{0}\sim\mu_{0}(\cdot\mid\phi),a_{0}\sim\pi(\cdot\mid s_{0},b_{\phi})}\left[Q_{R}^{\pi}(s_{0},a_{0},b_{\phi})+\alpha\cdot Q^{\pi}_{\text{safe}}(s_{0},a_{0},b_{\phi})\right]$
$=\mathbb{E}_{\phi\sim P_{\phi},s_{0}\sim\mu_{0}(\cdot\mid\phi),a_{0}\sim\pi(\cdot\mid s_{0},b_{\phi})}\left[Q_{R}^{\pi}(s_{0},a_{0},b_{\phi})+\alpha\cdot 0\right]$
$=J_{R}(\pi)$
Thus, maximizing $J_{\text{aug}}(\pi)$ within $\Pi_{\text{zero-violation}}$ is equivalent to maximizing $J_{R}(\pi)$ within the same set. ∎
@@
We define the safe state set for a given state $s\in S$, action $a\in A$, and hidden parameter $\phi\in\Phi$ as: $\mathcal{S}_{\mathrm{safe}}(s,a)=\left\{s^{\prime}\in S\mid C(s,a,s^{\prime})=0\text{ s.t. }T\left(s^{\prime}\mid s,a,\phi\right)>0\right\}.$ Thus, $\mathcal{S}_{\text{safe}}(s,a)$ contains next states $s^{\prime}$ where the safety condition is satisfied. Throughout the theorem, we address a cost function defined by safe distance